# Mamba2D: A Natively Multi-Dimensional State-Space Model for Vision Tasks

### v1

@@ Abstract
State-Space Models (SSMs) have recently emerged as a powerful and efficient alternative to the long-standing transformer architecture. However, existing SSM conceptualizations retain deeply rooted biases from their roots in natural language processing. This constrains their ability to appropriately model the spatially-dependent characteristics of visual inputs. In this paper, we address these limitations by re-deriving modern selective state-space techniques, starting from a natively multidimensional formulation. Currently, prior works attempt to apply natively 1D SSMs to 2D data (i.e. images) by relying on arbitrary combinations of 1D scan directions to capture spatial dependencies. In contrast, Mamba2D improves upon this with a single 2D scan direction that factors in both dimensions of the input natively, effectively modelling spatial dependencies when constructing hidden states. Mamba2D shows comparable performance to prior adaptations of SSMs for vision tasks, on standard image classification evaluations with the ImageNet-1K dataset.
@@ 2 Related Work
Moreover, the inherent causal properties of SSMs, combined with their sequential nature, imposes significant constraints on the flow of information between adjacent patches/pixels due to the 1-dimensional unrolling of their inputs. The recent works of V2M and 2DMamba aim to address these problems by introducing state transitions that account for dependencies in both horizontal and vertical directions, rather than adhering to a strictly linear sequence. Nevertheless, these methods still utilize independent sequential scans along each spatial axis. While this design enhances computational efficiency, it restricts the model’s capacity to effectively integrate information across neighbouring regions, thereby limiting its ability to fully exploit the spatial coherence inherent in image data.
@@ 4 Experiments
In this section, we present an experimental evaluation of the proposed framework. Firstly, we construct our M2D-T model with [3,3,9,3] blocks and [64,128,320,512] channels in each stage respectively, leading to a total model size of 26.7M parameters. Our model was trained on the ImageNet-1K dataset following the typical practices established in prior works [30, 18, 17]. Specifically, we employ a range of data augmentation schemes including: random cropping (to $224^{2}$), random horizontal flipping, auto random augmentation, MixUp and random erasing [36]. Whilst we also train for 300 epochs, we adopt the schedule free AdamW optimiser [3] with an effective batch size of 4096, a learning rate of 0.004, weight decay set to 0.05 and a 5% warmup w.r.t optimisation steps. Contrary to some contemporary works, we do not make use of EMA [28] as the schedule free optimiser acts as a more advanced form of weight averaging.

### v2

@@ Abstract
State-Space Models (SSMs) have recently emerged as a powerful and efficient alternative to the long-standing transformer architecture. However, existing SSM conceptualizations retain deeply rooted biases from their roots in natural language processing. This constrains their ability to appropriately model the spatially-dependent characteristics of visual inputs. In this paper, we address these limitations by re-deriving modern selective state-space techniques, starting from a natively multidimensional formulation. Currently, prior works attempt to apply natively 1D SSMs to 2D data (i.e. images) by relying on arbitrary combinations of 1D scan directions to capture spatial dependencies. In contrast, Mamba2D improves upon this with a single 2D scan direction that factors in both dimensions of the input natively, effectively modelling spatial dependencies when constructing hidden states. Mamba2D shows comparable performance to prior adaptations of SSMs for vision tasks, on standard image classification evaluations with the ImageNet-1K dataset. Source code is available at https://github.com/cocoalex00/Mamba2D.
@@ 2 Related Work
Moreover, the inherent causal properties of SSMs, combined with their sequential nature, imposes significant constraints on the flow of information between adjacent patches/pixels due to the 1-dimensional unrolling of their inputs. The recent works of V2M [32] and 2DMamba [37] aim to address these problems by introducing state transitions that account for dependencies in both horizontal and vertical directions, rather than adhering to a strictly linear sequence. Nevertheless, these methods still utilize independent sequential scans along each spatial axis. Similarly, Chimera [1] implements a 2D parallel scan for modelling multivariate time series data. Here, the full 2D spatial relationships are not required and a simplified transition matrix helps maintain efficiency. Although these designs enhance computational efficiency, the capacity of each method to effectively integrate information across neighbouring regions is restricted in various manners, thereby limiting the ability to fully exploit the spatial coherence inherent to image/vision data.
@@ 4 Experiments
In this section, we present an experimental evaluation of the proposed framework. Firstly, we construct our M2D-T model with [3,3,9,3] blocks and [64,128,320,512] channels in each stage respectively, leading to a total model size of 26.7M parameters. Our model was trained on the ImageNet-1K dataset following the typical practices established in prior works [31, 19, 18]. Specifically, we employ a range of data augmentation schemes including: random cropping (to $224^{2}$), random horizontal flipping, auto random augmentation, MixUp and random erasing [39]. Whilst we also train for 300 epochs, we adopt the schedule free AdamW optimiser [4] with an effective batch size of 4096, a learning rate of 0.004, weight decay set to 0.05 and a 5% warmup w.r.t optimisation steps. Additionally, inline with contemporary works, we also make use of EMA [29] during training.